OpenStudy (anonymous):

(3x-2)(2y-5) Is the Answer 60?

OpenStudy (mimi_x3):

No, where did you get 60 from ? Its, 6xy - 15x - 4y + 10

OpenStudy (anonymous):

I * 3 * 2 =6 and 2*5 = 10 then 6 * 10 = 60

OpenStudy (mimi_x3):

No, you don't times it , you expand its the simplify (: